OpenStudy (anonymous):

sin2x - sqrt(3)cos(2x) > sqrt(2) [Hint: use formula sin(a-b) = sin(a)cos(b) - cos(a)sin(b)]

OpenStudy (anonymous):

Hint2 sin(pi/3) = sqrt(3)/2

OpenStudy (anonymous):

In which order do I use these hints? I'm confused on how to break down the given equation so that I can begin to solve for the x values.

OpenStudy (anonymous):

try assuming that -sqrt(3)cos(2x) is actually -sin(y)cos(2x) using my hint, then use hint provided to merge them into 1

OpenStudy (anonymous):

I'm still a little lost sorry. What does the variable y represent and in your original hint, I don't see any component of it that I can apply (like where is sin(pi/3) or sqrt(3)/2 in the equation provided? ) Thanks replying by the way

OpenStudy (anonymous):

ok sin(pi/3) = sqrt(3)/2 as such sqrt(3) = 2sin(pi/3) therefore you can rewrite -sqrt(3)cos(2x) as -2sin(pi/3)cos(2x) ! Now it somewhat looks like a part of the hint doesnt it? =)

OpenStudy (anonymous):

Wow! Headsmack. I can't believe I missed that simple substitution that you were hinting at in your first reply. I'll give this a go again in a bit and make sure to award you Best Response after I've tried it (I assume that is the same as a medal?)

OpenStudy (anonymous):

good luck! I am still new to this this place so I have no idea what medals do/represent - but it should be easy from there, if you get stuck again - message away